### v1

@@ Abstract
on the unit ball of $\mathbb{R}^{N}$. We obtain the complete spectral structure of this problem. In particular, for $\alpha>0$, we find that the first eigenvalue is $k_{\nu,1}^{2}$ and the second eigenvalue is exactly $k_{\nu+1,1}^{2}$, where $k_{\nu+l,m}$ is the $m$th positive zero of $kJ_{\nu+l+1}(k)-(\alpha+l)J_{\nu+l}(k)$. Moreover, when $\alpha\in(-1,0)$, the first eigenvalue is $\widehat{k}_{\nu,1}^{2}$ where $\widehat{k}_{\nu,1}$ denotes the unique zero of $\alpha I_{\nu}(k)+kI_{\nu+1}(k)$, and the second eigenvalue is exactly $k_{\nu+1,1}^{2}$. Furthermore, for $\alpha=-1$, the first eigenvalue is $\widehat{k}_{\nu,1}^{2}$ and the second eigenvalue is exactly 0. Our conclusions indicate the ratio $\mu_{2}/\mu_{1}$ may be positive, negative or zero according to the suitable ranges of the parameter $\alpha$.
@@ 1 Introduction
where $\sigma\in\mathbb{R}$ is a physical constant and n is the unit outer normal vector to $\partial\Omega$. This condition indicates that the rate of change of temperature in the direction of the inner normal is proportional to the jump in temperature from the exterior to the interior of the body. For some given initial condition, by the method of separation of variables, one has the following Robin eigenvalue problem
@@
where $\beta\in\mathbb{R}$, which is also called the elastically supported membrane. For $N=1$, it is well known that problem (1.1) has and only has a sequence of simple eigenvalues. Moreover, the eigenfunctions corresponding to $k$th eigenvalue have k1k-1 simple zeros in $\Omega$ [7]. For $N=2,3$, when $\Omega$ is a ball, Courant and Hilbert [8] also gave the method for calculating eigenvalues and eigenfunctions to this problem but without specific calculation conclusion. For general domains or some special domains, a large number of mathematicians have studied this problem, and we will not list them one by one here, we refer to [12] and its references.
@@
The PPW conjecture has been included in problem lists of Yau [20, problem 77]. Yau commented that the truth of the conjecture means that one can determine whether a drum is circular or not by knowing its first two tones. The constant $3$ has been improved by several mathematicians. Specially, Brands [4] obtained $2.686$; then de Vries [18] obtained $2.658$; and finally, Chiti [6] obtained $2.586$. For any dimension, this conjecture can be stated that
@@
They proved that $\mu_{2}/\mu_{1}\leq 3$ for $\alpha\geq\alpha_{*}$ with some positive constant $\alpha_{*}$. In 2003, Henrot [11, Open problem 15] restated this conjecture by proposing the following open problem.
@@
Although the above question is only for a two-dimensional problem, it is evident that there are similar questions for high-dimensional problems, i.e., whether the ratio $\mu_{2}/\mu_{1}$ achieves its maximum for the ball for all values of $\alpha$ or for a range of values of $\alpha$. Different from the fixed membrane problem, we do not even know the exact value of $\mu_{2}/\mu_{1}$ for the ball. In fact, to the best of our knowledge, we even not know the exact value of $\mu_{2}/\mu_{1}$ in the one-dimensional case. So the main aim of this paper is to determine the exact value of $\mu_{2}/\mu_{1}$ on the ball, which may be the basis for solving the above conjecture or open problem.
@@
for $\alpha=-l$, the eigen-subspace corresponding to $\mu_{l,1}=0$ is generated by $r^{-\alpha}G_{-\alpha}(\xi)$; while, for $\alpha<-l$, the eigen-subspace corresponding to $\mu_{l,1}=-\widehat{k}_{\nu+l,1}^{2}<0$ is generated by $r^{-\nu}I_{\nu+l}\left(\widehat{k}_{\nu+l,1}r\right)G_{l}(\xi)$, where $G_{l}(\xi)$ is any eigenfunction corresponding to $\kappa_{l}=l(l+N-2)$ on N1N-1 dimensional sphere $\mathbb{S}^{N-1}$; in particular, for any $m\geq 2$, $\mu_{0,m}$ is simple and the corresponding eigenfunction is radially symmetric and has exactly m1m-1 simple zeros;
@@
for $\alpha>0$, the first eigenvalue is $k_{\nu,1}^{2}$ and the second eigenvalue is exactly $k_{\nu+1,1}^{2}$; when $\alpha\in(-1,0)$, the first eigenvalue is $\widehat{k}_{\nu,1}^{2}$ and the second eigenvalue is exactly $k_{\nu+1,1}^{2}$; for $\alpha=-1$, the first eigenvalue is $\widehat{k}_{\nu,1}^{2}$ and the second eigenvalue is exactly 0; moreover, $k_{\nu,1}$ and $k_{\nu+1,1}$ are strictly increasing with respect to $\alpha$, and as $\alpha$ tends to $\infty$, $k_{\nu,1}$ approaches $j_{\nu,1}$ and $k_{\nu+1,1}$ approaches $j_{\nu+1,1}$.
We mainly use the variable separation method (see, for example, [8]) to obtain the existence of eigenvalues and eigenfunctions. We further utilize some properties of Bessel functions to obtain the desired properties of eigenvalues and the corresponding eigenfunctions. In particular, we obtain the exact value of the second eigenvalue. Indeed, from Theorem 1.1 for $\alpha>0$ we know that
@@ 2 Spectrum of Robin problem in one dimension
It is well known [8, 9] that problem (2.1) has and only has a sequence of simple eigenvalues $\mu_{k}$ with $\mu_{1}<\mu_{2}<\cdots\rightarrow+\infty$. For $\mu\geq 0$, Bucur, Freitas and Kennedy [12, Chapter 4] gave the following relation of eigenvalue $\mu$ and the parameter $\alpha$
@@
From the monotonicity of $\alpha_{\pm}$ and its asymptotic behavior we can derive that $\mu_{1}$ and $\mu_{2}$ tend to -$\infty$ as $\alpha\rightarrow$-$\infty$. In particular, for $\alpha=-3$, from Proposition 2.1 we obtain that $\mu_{1}\approx-10.52118$ and $\mu_{2}\approx-6.63412$.
@@ 3 Spectrum of Robin problem on the ball
In this section, we establish the spectral structure of the Robin eigenvalue problem on the ball. For $N=2$ and $\mu\geq 0$, Bucur, Freitas and Kennedy [12, Chapter 4] derived the following transcendental relation between $\alpha$ and $\mu$
$\alpha J_{k}\left(\sqrt{\mu}\right)+\frac{1}{2}\sqrt{\mu}\left[J_{k-1}\left(\sqrt{\mu}\right)-J_{k+1}\left(\sqrt{\mu}\right)\right]=0$
@@
From the conclusions of [14, Theorem B ] or [10] we deduce that $\widetilde{h}_{\nu+l}(0)=-l$ and $\widetilde{h}_{\nu+l}(k)$ is strictly increasing in $\left(0,j_{\nu+l,1}\right)$, it has poles $j_{\nu+l,m}$ and is also strictly increasing between any two adjacent poles.
@@
From the conclusions of [14, Theorem C ] we deduce that $\widehat{h}_{\nu+l}(0)=-l$ and $\widehat{h}_{\nu+l}(k)$ is strictly decreasing in $\left(0,+\infty\right)$.
@@
up to a constant factor. In view of the boundary condition, one has that $l=-\alpha$.
@@
In particular, for any $m\geq 2$, the eigen-subspace corresponding to $\mu_{0,m}$ is one-dimensional which is generated by $v_{0,m}(r)$. So $\mu_{0,m}$ is simple and the corresponding eigenfunction $v_{0,m}(r)$ is radially symmetric. It has been known (see, for example, [5] or [7, Theorem 2.1 of Chap. 8]) that $v_{0,m}(r)$ has exactly m1m-1 simple zeros. So, we have proved property (b).
For any fixed $l\in\mathbb{N}\cup\{0\}$ and $\alpha\geq-l$, from the properties of $h_{\nu+l}(k)$, we derive that $k_{\nu+l,m}$ is strictly increasing with respect to $m$, which implies $\mu_{l,m}$ is strictly increasing with respect to $m$. Moreover, since $0\leq k_{\nu+l,1}<j_{\nu+l,1}$, $j_{\nu+l,m}<k_{\nu+l,m+1}<j_{\nu+l,m+1}$ and $j_{\nu+l,m}\rightarrow+\infty$ as $m\rightarrow+\infty$, we obtain that $k_{\nu+l,m}\rightarrow+\infty$ as $m\rightarrow+\infty$ which indicates that $\mu_{l,m}\rightarrow+\infty$ as $m\rightarrow+\infty$. Moreover, from [14, Theorem A] we know that $\widetilde{h}_{\nu+l}(k)$ is strictly decreasing with respect to $l$. Combining this with the monotonicity of $h_{\nu+l}(k)$ with respect to $k$ implies that $k_{\nu+l,m}$ with $\alpha\geq-l$ or $m\geq 2$ is strictly increasing with respect to $l$. Thus, when $\alpha\geq-l$ or $m\geq 2$, one sees that $\mu_{l,m}$ is strictly increasing with respect to $l$. So, property (c) is also verified.
@@
Moreover, it is clear from the monotonicity of $h_{\nu+l}$ that $k_{\nu,1}$ and $k_{\nu+1,1}$ are strictly increasing with respect to $\alpha$. As $\alpha$ tends to $\infty$, $k_{\nu,1}$ approaches $j_{\nu,1}$ and $k_{\nu+1,1}$ approaches $j_{\nu+1,1}$.
From the properties of ${h}_{\nu+l}(k)$ we can derive the following conclusions for $\alpha<0$. When $\alpha\in(-1,0)$, the first eigenvalue is $\widehat{k}_{\nu,1}^{2}$ where $\widehat{k}_{\nu,1}$ denotes the unique zero of $\alpha I_{\nu}(k)+kI_{\nu+1}(k)$, and the second eigenvalue is exactly $k_{\nu+1,1}^{2}$ where $k_{\nu+1,1}$ is the first positive zero of $kJ_{\nu+2}(k)-(\alpha+1)J_{\nu+1}(k)$. For $\alpha=-1$, the first eigenvalue is $\widehat{k}_{\nu,1}^{2}$ where $\widehat{k}_{\nu,1}$ denotes the unique zero of $I_{\nu}(k)+kI_{\nu+1}(k)$, and the second eigenvalue is exactly 0. For any given $n\in\mathbb{N}$ with $n\geq 2$, when $\alpha\in(-n,1-n)$, the first eigenvalue is $\min_{i\in\{0,\ldots,n-1\}}\left\{-\widehat{k}_{\nu+i,1}^{2}\right\}$ where $\widehat{k}_{\nu+i,1}$ denotes the unique zero of $\alpha I_{\nu+i}(k)+iI_{\nu+i}(k)+kI_{\nu+i+1}(k)$. For $n\in\mathbb{N}$ with $n\geq 2$ and $\alpha=-n$, the first eigenvalue is $\min_{i\in\{0,\ldots,n-1\}}\left\{-\widehat{k}_{\nu+i,1}^{2}\right\}$ where $\widehat{k}_{\nu+i,1}$ denotes the unique zero of $nI_{\nu+i}(k)+iI_{\nu+i}(k)+kI_{\nu+i+1}(k)$. So we obtain the desired conclusions of (d). ∎
From the argument of Theorem 1.1, we also obtain $n$ negative eigenvalues $\widehat{k}_{\nu+i,1}^{2},i\in\{0,\ldots,n-1\}$ when $\alpha\in(-n,1-n)$ with $2\leq n\in\mathbb{N}$. While, for $\alpha=-n$ with $2\leq n\in\mathbb{N}$, besides $n$ negative eigenvalues, 0 is also an eigenvalue. Undoubtedly, the smallest among these $n$ negative eigenvalues is the first eigenvalue, followed by the second eigenvalue. But since we cannot accurately sort them by size, we cannot determine who is the first two, which is fundamentally different from one-dimensional problem (see Proposition 2.1). We conjecture that $\widehat{k}_{\nu+i,1}^{2}$ is decreasing with respect to $i$. If this conjecture was valid, the first eigenvalue is just $\widehat{k}_{\nu,1}^{2}$ and the second eigenvalue is $\widehat{k}_{\nu+1,1}^{2}$. Note that this conjecture can be derived from the monotonicity of $\widehat{h}_{\nu+l}(k)$ with respect to $l$. Thus, we also conjecture that $\widehat{h}_{\nu+l}(k)$ is decreasing with respect to $l$.
Since $I_{\nu+1}(k)/I_{\nu}(k)$ increases to $1$ as $k$ increases to $\infty$, we find that $\widehat{h}_{\nu+l}(k)\rightarrow$-$\infty$ (monotonously) as $k\rightarrow+\infty$. This implies that $\mu_{l,1}\rightarrow$-$\infty$ as $\alpha\rightarrow$-$\infty$. In particular, the first eigenvalue tends (monotonously) to -$\infty$ as $\alpha\rightarrow$-$\infty$.

### v2

@@ Abstract
on the unit ball of $\mathbb{R}^{N}$. We obtain the complete spectral structure of this problem. In particular, for $\alpha>0$, the first eigenvalue is $k_{\nu,1}^{2}$ and the second eigenvalue is $k_{\nu+1,1}^{2}$, where $k_{\nu+l,m}$ is the $m$th positive zero of $kJ_{\nu+l+1}(k)-(\alpha+l)J_{\nu+l}(k)$. Moreover, when $\alpha\in(-l,1-l)$ with any $l\in\mathbb{N}$, one has $l$ negative (strictly increasing) eigenvalues $\widehat{k}_{\nu+i,1}^{2}$ with $i\in\{0,\ldots,l-1\}$ where $\widehat{k}_{\nu+l,1}$ denotes the unique zero of $\alpha I_{\nu+l}(k)+lI_{\nu+l}(k)+kI_{\nu+l+1}(k)$; while, for $\alpha=-l$, besides $l$ negative (increasing) eigenvalues, 0 is also an eigenvalue.
@@ 1 Introduction
where $\sigma$ is a positive physical constant and n is the unit outer normal vector to $\partial\Omega$. This condition indicates that the rate of change of temperature in the direction of the inner normal is proportional to the jump in temperature from the exterior to the interior of the body. For some given initial condition, by the method of separation of variables, one has the following Robin eigenvalue problem
@@
where $\beta\in\mathbb{R}$, which is also called the elastically supported membrane. This problem with positive boundary parameter is also related to the wave equation to elastically restoring boundaries. Negative Robin parameter is arising in a model for surface superconductivity [15]. For $N=1$, it is well known that problem (1.1) has and only has a sequence of simple eigenvalues. Moreover, the eigenfunctions corresponding to $k$th eigenvalue have k1k-1 simple zeros in $\Omega$ [8]. For $N=2,3$, when $\Omega$ is a ball, Courant and Hilbert [9] also gave the method for calculating eigenvalues and eigenfunctions to this problem but without specific calculation conclusion. For general domains or some special domains, a large number of mathematicians have studied this problem, and we will not list them one by one here, we refer to [19] and its references.
@@
The PPW conjecture has been included in problem lists of Yau [29, problem 77]. Yau commented that the truth of the conjecture means that one can determine whether a drum is circular or not by knowing its first two tones. The constant $3$ has been improved by several mathematicians. Specially, Brands [5] obtained $2.686$; then de Vries [27] obtained $2.658$; and finally, Chiti [7] obtained $2.586$. For any dimension, this conjecture can be stated that
@@
They proved that $\mu_{2}/\mu_{1}\leq 3$ for $\alpha\geq\alpha_{*}$ with some positive constant $\alpha_{*}$. In 2003, Henrot [18, Open problem 15] restated this conjecture by proposing the following open problem.
@@
Although the above question is only for a two-dimensional problem, it is evident that there are similar questions for high-dimensional problems, i.e., whether the ratio $\mu_{2}/\mu_{1}$ achieves its maximum for the ball for all values of $\alpha$ or for a range of values of $\alpha$.
Recently, Freitas and Laugesen [13] investigated the first and second eigenvalues for the ball (also see [14] for $2$-dimensional case). Then Langford and Laugesen [21] treated the first and second Robin eigenvalues for geodesic disks in Euclidean, hyperbolic and spherical spaces, which does not use special functions. When $\alpha>0$, Freitas [10] obtained sharp estimates for the first eigenvalue of the ball and related quantities. While, for negative boundary parameter, Antunes, Freitas and Krejčiřík [1] and Freitas and Krejčiřík [12] obtained some bounds for $\mu_{1}$. In [22], Laugesen investigated various optimization problems of the first two eigenvalues on rectangles.
Different from the fixed membrane problem, the value of $\mu_{2}/\mu_{1}$ for the ball is dependent on parameter $\alpha$. Here, based on the above important conclusions, we plan to obtain the complete spectrum of the Robin eigenvalue problem on the ball. In particular, determine the exact value of $\mu_{2}/\mu_{1}$ on the ball, which may be the basis for solving the above conjecture or open problem.
@@
for $\alpha=-l$, the eigen-subspace corresponding to $\mu_{l,1}=0$ is generated by $r^{-\alpha}G_{-\alpha}(\xi)$; while, for $\alpha<-l$, the eigen-subspace corresponding to $\mu_{l,1}=-\widehat{k}_{\nu+l,1}^{2}<0$ is generated by $r^{-\nu}I_{\nu+l}\left(\widehat{k}_{\nu+l,1}r\right)G_{l}(\xi)$, where $G_{l}(\xi)$ is any eigenfunction corresponding to $\kappa_{l}=l(l+N-2)$ on N1N-1 dimensional sphere $\mathbb{S}^{N-1}$; in particular, $\mu_{0,1}$ is simple with positive radial eigenfunction, for any $m\geq 2$, $\mu_{0,m}$ is simple and the corresponding eigenfunction is radially symmetric and has exactly m1m-1 simple zeros;
@@
for $\alpha>0$, the first eigenvalue is $k_{\nu,1}^{2}$ and the second eigenvalue is exactly $k_{\nu+1,1}^{2}$. Moreover, $k_{\nu,1}$ and $k_{\nu+1,1}$ are strictly increasing with respect to $\alpha$, and as $\alpha$ tends to $\infty$, $k_{\nu,1}$ approaches $j_{\nu,1}$ and $k_{\nu+1,1}$ approaches $j_{\nu+1,1}$;
(e) when $\alpha\in(-l,1-l)$ with any $l\in\mathbb{N}$, one has $l$ negative (strictly increasing) eigenvalues $\widehat{k}_{\nu+i,1}^{2},i\in\{0,\ldots,l-1\}$; while, for $\alpha=-l$ with any $l\in\mathbb{N}$, besides $l$ negative (increasing) eigenvalues, 0 is also an eigenvalue. In particular, when $\alpha\in(-1,0)$, the first eigenvalue is $\widehat{k}_{\nu,1}^{2}$ and the second eigenvalue is exactly $k_{\nu+1,1}^{2}$; for $\alpha=-1$, the first eigenvalue is $\widehat{k}_{\nu,1}^{2}$ and the second eigenvalue is exactly 0.
As mentioned earlier, the existence, and upper and lower bounds estimation of the first two eigenvalues have been obtained in references [1, 13, 12, 10, 14, 21, 22]. Here the complete spectrum is been obtained, and the precise calculation formulas (through Bessel functions) for the first two are provided. Moreover, we can also determine the exact number of negative eigenvalues based on the range of boundary parameter. These conclusions can serve as a supplement to the important literature mentioned above and also as a basis for further research on related problems. It should be noted that some of the conclusions regarding the first two eigenvalues may have been covered by the aforementioned literatures, but for the sake of completeness, we still list them here.
We mainly use the variable separation method (see, for example, [9] or [13]) to obtain the existence of eigenvalues and eigenfunctions. We further utilize some properties of Bessel functions to obtain the desired properties of eigenvalues and the corresponding eigenfunctions. In particular, we obtain the calculation formula for the first two eigenvalues. Indeed, from Theorem 1.1 for $\alpha>0$ we know that
@@
In [22, Conjecture B], Laugesen conjectured that $\mu_{2}/\mu_{1}$ on bounded Lipschitz domain is decreasing for $\alpha>0$. The data in Tables 1 and 2 supports this conjecture (at least on the ball).
@@ 2 Spectrum of Robin problem in one dimension
It is well known [9, 16] that problem (2.1) has and only has a sequence of simple eigenvalues $\mu_{k}$ with $\mu_{1}<\mu_{2}<\cdots\rightarrow+\infty$. For $\mu\geq 0$, Bucur, Freitas and Kennedy [19, Chapter 4] gave the following relation of eigenvalue $\mu$ and the parameter $\alpha$
@@
As far as we know, the key value $\alpha=-2$ is our new discovery. This point is the key point for the change in the number of negative eigenvalues. When $\alpha<-2$, there are exactly two negative eigenvalues, and when $\alpha\in[-2,0)$, there is exactly one negative eigenvalue. From the monotonicity of $\alpha_{\pm}$ and its asymptotic behavior we can derive that $\mu_{1}$ and $\mu_{2}$ tend to -$\infty$ as $\alpha\rightarrow$-$\infty$. In particular, for $\alpha=-3$, from Proposition 2.1 we obtain that $\mu_{1}\approx-10.52118$ and $\mu_{2}\approx-6.63412$. For $\alpha>0$, we have that $\mu_{1}\in(0,\pi)$ and $\mu_{2}\in(\pi,2\pi)$. In this case, Freitas and J.B. Kennedy [11] obtained more refined upper and lower bound estimates for the first two eigenvalues.
@@ 3 Spectrum of Robin problem on the ball
In this section, we establish the spectral structure of the Robin eigenvalue problem on the ball. For $N=2$ and $\mu\geq 0$, Bucur, Freitas and Kennedy [19, Chapter 4] derived the following transcendental relation between $\alpha$ and $\mu$
$\alpha J_{k}\left(\sqrt{\mu}\right)+\frac{1}{2}\sqrt{\mu}\left[J_{k-1}\left(\sqrt{\mu}\right)-J_{k+1}\left(\sqrt{\mu}\right)\right]=0,$
@@
The above equation also derived by Freitas [10, Section 3] or [13, Section 5] for the special case of $l=0$ and $\alpha>0$. From the conclusions of [23, Theorem B ] or [17] we deduce that $\widetilde{h}_{\nu+l}(0)=-l$ and $\widetilde{h}_{\nu+l}(k)$ is strictly increasing in $\left(0,j_{\nu+l,1}\right)$, it has poles $j_{\nu+l,m}$ and is also strictly increasing between any two adjacent poles.
@@
From the conclusions of [23, Theorem C ] we deduce that $\widehat{h}_{\nu+l}(0)=-l$ and $\widehat{h}_{\nu+l}(k)$ is strictly decreasing in $\left(0,+\infty\right)$.
@@
up to a constant factor. In view of the boundary condition, one has that $l=-\alpha$, which also have been obtained in [13, Section 5].
@@
In particular, $\mu_{0,1}$ is simple with positive radial eigenfunction, and for any $m\geq 2$, the eigen-subspace corresponding to $\mu_{0,m}$ is one-dimensional which is generated by $v_{0,m}(r)$. So $\mu_{0,m}$ is simple and the corresponding eigenfunction $v_{0,m}(r)$ is radially symmetric. It has been known (see, for example, [6] or [8, Theorem 2.1 of Chap. 8]) that $v_{0,m}(r)$ has exactly m1m-1 simple zeros. So, we have proved property (b).
For any fixed $l\in\mathbb{N}\cup\{0\}$ and $\alpha\geq-l$, from the properties of $h_{\nu+l}(k)$, we derive that $k_{\nu+l,m}$ is strictly increasing with respect to $m$, which implies $\mu_{l,m}$ is strictly increasing with respect to $m$. Moreover, since $0\leq k_{\nu+l,1}<j_{\nu+l,1}$, $j_{\nu+l,m}<k_{\nu+l,m+1}<j_{\nu+l,m+1}$ and $j_{\nu+l,m}\rightarrow+\infty$ as $m\rightarrow+\infty$, we obtain that $k_{\nu+l,m}\rightarrow+\infty$ as $m\rightarrow+\infty$ which indicates that $\mu_{l,m}\rightarrow+\infty$ as $m\rightarrow+\infty$. Moreover, from [23, Theorem A] we know that $\widetilde{h}_{\nu+l}(k)$ is strictly decreasing with respect to $l$. Combining this with the monotonicity of $h_{\nu+l}(k)$ with respect to $k$ implies that $k_{\nu+l,m}$ with $\alpha\geq-l$ or $m\geq 2$ is strictly increasing with respect to $l$. Thus, when $\alpha\geq-l$ or $m\geq 2$, one sees that $\mu_{l,m}$ is strictly increasing with respect to $l$. So, property (c) is also verified.
@@
Moreover, it is clear from the monotonicity of $h_{\nu+l}$ that $k_{\nu,1}$ and $k_{\nu+1,1}$ are strictly increasing with respect to $\alpha$. As $\alpha$ tends to $\infty$, $k_{\nu,1}$ approaches $j_{\nu,1}$ and $k_{\nu+1,1}$ approaches $j_{\nu+1,1}$. So we obtain the desired conclusions of (d).
Finally we show property (e). From the properties of ${h}_{\nu+l}(k)$ we can derive the following conclusions for $\alpha<0$. When $\alpha\in(-1,0)$, the first eigenvalue is $\widehat{k}_{\nu,1}^{2}$ where $\widehat{k}_{\nu,1}$ denotes the unique zero of $\alpha I_{\nu}(k)+kI_{\nu+1}(k)$, and the second eigenvalue is exactly $k_{\nu+1,1}^{2}$ where $k_{\nu+1,1}$ is the first positive zero of $kJ_{\nu+2}(k)-(\alpha+1)J_{\nu+1}(k)$. For $\alpha=-1$, the first eigenvalue is $\widehat{k}_{\nu,1}^{2}$ where $\widehat{k}_{\nu,1}$ denotes the unique zero of $I_{\nu}(k)+kI_{\nu+1}(k)$, and the second eigenvalue is exactly 0. In general, for any given $l\in\mathbb{N}$ with $l\geq 2$, when $\alpha\in(-l,1-l)$, the first eigenvalue is $\min_{i\in\{0,\ldots,l-1\}}\left\{-\widehat{k}_{\nu+i,1}^{2}\right\}$ where $\widehat{k}_{\nu+i,1}$ denotes the unique zero of $\alpha I_{\nu+i}(k)+iI_{\nu+i}(k)+kI_{\nu+i+1}(k)$. While, for $l\in\mathbb{N}$ with $l\geq 2$ and $\alpha=-l$, the first eigenvalue is $\min_{i\in\{0,\ldots,l-1\}}\left\{-\widehat{k}_{\nu+i,1}^{2}\right\}$ where $\widehat{k}_{\nu+i,1}$ denotes the unique zero of $lI_{\nu+i}(k)+iI_{\nu+i}(k)+kI_{\nu+i+1}(k)$. From [13, Lemma 10] we derive that $\widehat{h}_{\nu+l}(k)$ is strictly decreasing with respect to $l$. This implies that $\widehat{k}_{\nu+i,1}^{2}$ is strictly decreasing with respect to $i$. Hence, $\widehat{k}_{\nu+i,1}^{2}$ is strictly increasing with respect to $i$. In particular, the first eigenvalue is just $\widehat{k}_{\nu,1}^{2}$ and the second eigenvalue is $\widehat{k}_{\nu+1,1}^{2}$. ∎
From Theorem 1.1 we can see that there are infinite negative eigenvalues when $\alpha\rightarrow$-$\infty$, which is fundamentally different from one-dimensional problem (see Proposition 2.1). Since $I_{\nu+1}(k)/I_{\nu}(k)$ increases to $1$ as $k$ increases to $\infty$, we find that $\widehat{h}_{\nu+l}(k)\rightarrow$-$\infty$ (monotonously) as $k\rightarrow+\infty$. This implies that $\mu_{l,1}\rightarrow$-$\infty$ as $\alpha\rightarrow$-$\infty$. In particular, the first eigenvalue tends (monotonously) to -$\infty$ as $\alpha\rightarrow$-$\infty$.